OpenStudy (anonymous):

Simplify. Express the quotient as a radical expression. sqrt[9]{x^7} / sqrt[9]{x^5}

OpenStudy (anonymous):

Is it?? \[\huge \frac{\sqrt{9}x^7}{\sqrt{9}x^5}\]

OpenStudy (anonymous):

OR?? \[\huge \frac{\sqrt{9x^7}}{\sqrt{9x^5}}\]

OpenStudy (anonymous):

\[\sqrt[9]{x^7} =x^{7/9}\] \[\sqrt[9]{x^5} =x^{5/9}\] Now \[\sqrt[9]{x^7} / \sqrt[9]{x^5} = x^{7/9}/x^{5/9}\] \[= x^{7/9 - 5/9}\] \[= x^{2/9}\] or \[\sqrt[9]{x^{2}}\]